OpenStudy (anonymous):

Prove that: \[1.\frac {tanA+secA-1}{tanA-secA+1}=\frac{1+sinA}{cosA}\] and \[2.(tanA+cosecB)^{2}-(cotB-secA)^{2}=2tanAcotB(cosecA+secB)\]

OpenStudy (anonymous):

You can substitute: \[1 = sec^2 \theta - tan^2 \theta\]

OpenStudy (anonymous):

In the first one...

OpenStudy (anonymous):

full sol. pls

OpenStudy (anonymous):

Or you can say: \[1 = sec^2 \theta - tan^2 \theta = (sec \theta + tan \theta)(sec \theta - tan \theta)\] Try to substitute it there and show me what you got..

OpenStudy (anonymous):

i will do it on paper.. wait a moment

OpenStudy (anonymous):

\[\frac{(\tan A + secA) (tanA - secA +1)}{(tanA - secA +1)} \implies tanA + \sec A\]

OpenStudy (anonymous):

As waterineyes said, put 1 = (secA+tanA)(secA-tanA) in denominator.. we will evaluate denominator first. D = tanA-secA+(tanA+secA)(secA-tanA) take (secA-tanA) common (secA-tanA)[tanA+secA-1] now numerator = tanA+secA-1.. numerator and denominator will cancel tanA+secA-1.. which leaves expression as 1 --------- ... secA-tanA we know that (secA+tanA)(secA-tanA)=1 => secA+tanA = 1/(secA-tanA) this makes the above expression = secA+tanA =(1/cosA) + (sinA/cosA) =(1+sinA)/cosA I tried my best lol.. :) it takes too much time to type.. :P so try 2nd one yourself

OpenStudy (anonymous):

\[tanA + secA = \frac{sinA}{cosA} + \frac{1}{cosA} = \frac{1 + sinA}{cosA}\]

OpenStudy (anonymous):

lol ^^ nice one :P

OpenStudy (anonymous):

should have done this

OpenStudy (anonymous):

@shubham.bagrecha u there bro ?

OpenStudy (anonymous):

\[\frac{tanA + secA - (tanA + secA)(tanA - secA)}{tanA - secA + 1}\] \[\frac{(tanA + secA)(1 - secA + \tan A)}{\tan A - \sec A + 1} \implies \frac{(tanA + secA)(\tan A - \sec A + 1)}{\tan A - \sec A + 1}\] \[\implies tanA + secA \implies \frac{1 + sinA}{cosA}\]

OpenStudy (anonymous):

He is only here for full solutions..

OpenStudy (anonymous):

I remember there was same question like this in my final exams :P

OpenStudy (anonymous):

now 2nd ques.

OpenStudy (anonymous):

More precisely if I can guess It was in CBSE 2003 In India..

OpenStudy (anonymous):

I gave my 10th standard paper in March 2012 lol

OpenStudy (anonymous):

anyways i am attempting 2nd one

OpenStudy (anonymous):

Ha ha ha..

OpenStudy (anonymous):

hey i am stuck .. hehe

OpenStudy (anonymous):

Very nice question.. Love to do the second one..

OpenStudy (anonymous):

In second the identities using are: \[\sec^2A - \tan^2 A = 1\] \[cosec^2 A - \cot^2 A = 1\] \[cotB = \frac{cosB}{sinB}\] \[tanA = \frac{sinA}{cosA}\]

OpenStudy (anonymous):

So, firstly open the brackets by applying the squares to the brackets: \[\tan^2A + cosec^2 A + 2tanA \cdot cosecA - \cot^2B - \sec^2B +2cotB\cdot secB\] After applying the first two identities: 1 and -1 will cancel out: \[\implies 2(tanA cosecA + cotBsecB)\] Take \(tanAcotB\) common: \[\implies 2tanA cotB(\frac{cosecB}{cotB} + \frac{secA}{tanA})\] \[\implies 2tanA cotB(cosecA + secB)\]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Second last step I simplified as: \[\frac{cosecB}{cotB} + \frac{secA}{tanA} \implies \frac{cosecB \cdot sinB}{cosB} + \frac{secA \cdot cosA}{sinA}\] \[\implies \frac{1}{cosB} + \frac{1}{sinA} \implies (secB + cosecA)\]

OpenStudy (anonymous):

Welcome dear..